OpenStudy (anonymous):

plz help me in ques no.2 part 1

OpenStudy (anonymous):

@axi Vx = V*(R2/(R1+R2)-R4/(R3+R4)) R? = (R2*R3)/R1

OpenStudy (anonymous):

thnx and part 2 ?

OpenStudy (anonymous):

tistaroun have solved q5?/

OpenStudy (anonymous):

@axi Vth = V * ( R1/(R1+R3) - R2/(R2+R4) ) Rth = R1*R2/(R1+R2) + R3*R4/(R3+R4) Vth modified the same as Vth Rth modified = Rth/5

OpenStudy (anonymous):

plz tistaroun plz help in Q5

OpenStudy (anonymous):

tnx @tistaroun

OpenStudy (anonymous):

@tistaroun part 3 plzplz i m getting wrong answer nw only one attempt is leftt plz help me out

OpenStudy (anonymous):

@axi iR,max = Vs/R iR,min = 0

OpenStudy (anonymous):

tistaroun ????? q5

OpenStudy (anonymous):

@tistaroun only one ques is left plz help me out with q6 plzzz

OpenStudy (anonymous):

@axi what do you need?

OpenStudy (anonymous):

its answers plzz

OpenStudy (anonymous):

@axi But you may have another values.

OpenStudy (anonymous):

Here we examine noise in a digital channel between an AND gate and an inverter shown in Figure 16. The channel has a length of D=26m and noise is added in at the rate of 0.6Vm. Figure 16 To correct for that noise, we introduce non-inverting buffers into the channel. The purpose of a buffer is to take an input and replicate it at the buffer's output. This buffer is useful because it clean up a signal which no longer meets the ouptut specification of the static discipline due to noise, so that it once again meets the output specifications of the static discipline. The buffered channel is show in Figure 17 below. Figure 17 These buffers, as well as the gates shown, obey the following static discipline: VOL=1V,VIL=2V,VIH=3V,VOH=4V. What is the minimum number of buffers required to connect between the digital links in order to ensure correct operation? unanswered Now let's consider the NOR gate in Figure 18 which has VS=5V and the following given static discipline: VOL=0.5V,VIL=1V,VIH=4V,VOH=4.6V. Figure 18 Assume that both MOSFETs have the same VT and RON=10kΩ. What is the minimum value of VT in Volts that allows the NOR gate to satisfy the given static discipline? unanswered What is the maximum value of VT in Volts that allows the NOR gate to satisfy the given static discipline? unanswered What is the minimum value of RL (in kΩ) that allows the NOR gate to satisfy the given static discipline? unanswered You have used 0 of 3 submissions Q6: PART 2 Consider the circuit in Figure 19 below. Figure 19 Write a boolean expression for Z in terms of A,B,C,D, and E. You need not simplify your expression. Use your expression to fill out the truth table below. A B C D E Z --- --- --- --- --- --- 1 1 0 0 1 z0 1 1 1 0 1 z1 0 0 1 1 1 z2 0 1 1 0 1 z3 Enter the unknown values for the ouputs as one stream of bits in the form z0z1z2z3 unanswered How many distinct boolean-valued functions are there of n boolean-valued signals? Write an expression in terms of n. unanswered

OpenStudy (anonymous):

plzz help me Q5 answer tell me........plzzz

OpenStudy (anonymous):

Q6 Part 2 1011 2^(2^n)

OpenStudy (anonymous):

Q6 Part1 buffers = (D*noise rate)/(Vil - Vol) minimum value of VT = Vil maximum value of VT = Vih minimum value of RL = ((Vs*Ron)/Vol) - Ron

OpenStudy (anonymous):

@axi Now you should close this question.

OpenStudy (anonymous):

kk thnx A lot @tistaroun

OpenStudy (anonymous):

@axi You are welcome.

OpenStudy (anonymous):

ANYONE have equations for Q5 plz

OpenStudy (anonymous):

A B C D E Z --- --- --- --- --- --- 1 0 1 0 1 z0 1 1 0 1 1 z1 0 0 1 1 0 z2 0 0 1 0 0 z3 For this Input the value of z0z1z2z3 is?